OpenStudy (anonymous):

Find the magnitude ||v|| and direction angle θ, to the nearest tenth of a degree, for the given vector v. v = -4i - 3j

OpenStudy (anonymous):

if w= a i + b j \[ || w|| = \sqrt{ a^2 + b^2} \] What is the norm of u=-4i - 3j ?

OpenStudy (anonymous):

\[ ||u|| = \sqrt{ (-4)^2 + (-3)^2}=\sqrt{ 25}=5 \]

OpenStudy (anonymous):

So would this 216.9° be the other part?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

YW